An asset used in a 4-year project falls in the 5-year MACRS class for tax purposes. The asset has an acquisition cost of $8,300,000 and will be sold for $1,960,000 at the end of the project. If the tax rate is 22 percent, what is the aftertax salvage value of the asset? Refer to (MACRS schedule) (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567)
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Acquisition cost         8,300,000.00
Depreciation for 4 years in % = 20% + 32% + 19.20%+11.52% 82.72%
Depreciation in $         6,865,760.00
Carrying amount at end of 4 years = 8300,000 - 6865,760         1,434,240.00
Sale Value         1,960,000.00
Gain on Sale             525,760.00
Tax on Gain = 525,760 *22%             115,667.20
aftertax salvage value of the asset = 1960,000 - 115,667.20         1,844,332.80